### v1

# CosForce: A Force-Based General Model for Simulating Pedestrian Anticipation and Reaction Mechanisms

Abstract
In this study, a force-based general model, named the CosForce model, has been developed. To the best of our knowledge, this may represent the most simplified version of the social force model (streamlined rules and high efficiency). Considering the anisotropic influence of the environment on pedestrian, the model solely focuses on interactions between pedestrians and their nearest neighbors, rather than all surrounding pedestrians (as the pedestrian count increases, the computational cost transitions from power-law growth to linear growth). Specifically, cosine functions are employed to describe the interactions between nearest neighbors, allowing the model to simulate the anticipation and reaction mechanisms. By applying this model to simulations across various scenarios, a wide range of pedestrian dynamics can be simulated, including pedestrian fundamental diagrams and self-organized phenomena such as lane formation, stripe formation, and cross-channel formation, among others.
The open-source codes of the model can be found at: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link, and a short video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.
Force-based model, Pedestrian self-organization, Collective dynamics, Numerical simulation
1 Introduction
The study of pedestrian dynamics modeling has undergone several decades of development. As a branch of collective dynamics, it has naturally integrated with fields such as multi-agent systems and robotic planning, currently exhibiting unprecedented vitality. Microscopic pedestrian dynamics models primarily focus on crowd simulation driven by pedestrian interactions. The fundamental models generally consider only the real-time interaction processes between pedestrians and the environment. On a broader scale, additional rule sets can be incorporated to simulate mechanisms such as route decision in pedestrian traffic, etc. Classic approaches, such as the cellular automata model (Burstedde et al., 2001), social force model (Helbing & Molnar, 1995), reciprocal velocity obstacles model (Van den Berg et al., 2008) and the heuristic model (Moussaïd et al., 2011), remain mainstream methods.
Due to their high extensibility, force-based models have been applied not only in pedestrian dynamics modeling but also in broader fields (Liu et al., 2024). New rule sets can be added by researchers to investigate specific mechanisms as required. Among these mechanisms, pedestrian anticipation and reaction have been extensively analyzed. Pedestrian reaction behaviors are generally viewed as passive responses to the environment, such as car-following behavior, whereas anticipation behaviors are considered active responses to environmental stimuli, such as a goalkeeper’s predictive dive in soccer. The subtle distinctions between anticipation and reaction mechanisms are frequently overlooked, leading to the homogenization of the modeling process. Collision avoidance rules can be introduced into force-based models in various forms, based on the relative velocity and relative distance of agents. Currently, the social force model incorporating collision avoidance rules based on the time-to-collision (TTC) estimation is commonly regarded as a model embedding anticipation mechanisms (Zanlungo et al., 2011; Lü et al., 2020; Xu et al., 2021; Liu et al., 2024).
A potential issue is that the addition of rule sets or new force terms may render the model cumbersome, thereby reducing its generalizability. Such models are often effective in specific scenarios, but their optimization for specific objectives may, in fact, weaken their generalization capability. In this regard, it may be more effective to develop a general model by simplifying existing models rather than by adding complexity. To this end, a model was developed to simulate anticipation and reaction mechanisms in pedestrian interaction. The rules of the model were further simplified based on the classic social force model while aiming for higher simulation accuracy.
2 Experience and Intuition in Pedestrian Interactions
Pedestrians resemble elastic particles rather than rigid particles
In the classical social force model, pedestrians are represented as rigid particles. The volume exclusion principle, which is strictly constrained by the function $g(x)$, has been shown to potentially cause distortions in pedestrian dynamics under extreme conditions. In order to capture more detailed motion dynamics, some studies have proposed the use of ellipsoidal shapes instead of circular ones to simulate pedestrian movement (Chraibi et al., 2010). However, this modification inevitably results in increased model complexity. A more effective compromise can be achieved by considering pedestrians as elastic bodies rather than rigid ones. In narrow passages, pedestrians frequently turn sideways, and mutual compression between individuals reduces the space occupied. Therefore, modeling pedestrians as elastic bodies allows for a more accurate simulation of these behaviors.
Figure 1: Scaling down the laws of stripe formation, (a) schematic diagram of stripe formation, (b) decision payoff matrix for pedestrians in binary conflicts, (c) cooperative and non-cooperative games between conflicting pedestrians.
Anisotropy of pedestrian motion space
Traditionally, the space surrounding pedestrians has been assumed to be isotropic, implying that the influence exerted by pedestrians in different directions on a target pedestrian is identical. The interaction force between pedestrians has typically been modeled as an exponentially decaying function of distance. The rapid decay property ensures that force interactions are significant only between nearby pedestrians. However, this modeling approach results in a potential issue: forces from various directions may produce opposing effects, causing pedestrians to maintain high speeds even in high-density situations. Such results contradict empirical observations of pedestrian fundamental diagrams (Parisi et al., 2009).
In real-world movement, it is generally observed that pedestrians primarily focus on the dynamics ahead of them while largely disregarding those behind them. Therefore, incorporating anisotropy in modeling pedestrian movement appears to be more consistent with the intuitive and empirically observation of pedestrians.
Cooperative game induces local self-organization
The phenomenon of self-organization among pedestrians is characterized by entropy reduction and the emergence of steady states within pedestrian flow systems. We conjecture that such entropy reduction is presumably induced by local cooperative interactions among conflicting pedestrians (Bonnemain et al., 2023). For instance, in a typical crossing flow scenario, as depicted in Fig. 1(a), there is a potential risk of collision between crossing pedestrians. Anticipation mechanism, involving a negotiation or game process, can facilitate phase separation between pedestrians moving in different directions. The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig. 1(b). Cooperative games contribute to entropy reduction within the system, leading to the emergence of spontaneous order, as illustrated in Fig. 1(c).
Collisions are necessary
The development of models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016). However, a potential issue with this model is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable. This discrepancy arises mainly because pedestrian speed and movement direction are updated based on empirical estimates rather than precise numerical calculations. In reality, pedestrians can’t perform such calculations, so their movements rely on rough approximations. In this consideration, the collision process is crucial for pedestrian simulation, especially in high-density crowd scenarios.
3 Model
3.1 Model Rules
Prior to introducing the model, the Horizontal Field of Attention (HFA, $\boldsymbol{\Omega}$) is defined. The HFA is constrained by the eccentricity attentional angle $\phi$ and the attentional depth $h$, as depicted in Fig. 2. The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement. The attentional depth $h$ can theoretically extend to infinity. For the purpose of minimizing computational cost, $h$ can be restricted to the minimum headway that allows pedestrians to maintain free motion.
Figure 2: Diagram of the process of pedestrian interaction.
In analogy to the classical Social Force Model, the pedestrian movement in the proposed model is described as a combination of a self-driving force, an inter-pedestrian interaction force, and an obstacle-pedestrian interaction force. With respect to the inter-pedestrian interaction force, due to the anisotropy of the motion space, only the force $f_{ij}$, exerted by the nearest pedestrian $j$ within the region $\boldsymbol{\Omega}$, is considered. The global constraints can be expressed as follows:
$\boldsymbol{f}={{\boldsymbol{f}}_{i}}+{{\boldsymbol{f}}_{ij}}+\sum\limits_{w}{{{\boldsymbol{f}}_{iw}}}$ (1)
Here, $\boldsymbol{f}$ represents the net force acting on a target pedestrian, $\boldsymbol{f}_{i}$ denotes the self-driving force of the pedestrian, $\boldsymbol{f}_{ij}$ represents the interaction force from the nearest neighbor $j$, and $\boldsymbol{f}_{iw}$ represents the force exerted by obstacles such as walls.
The self-driving force is driven by the free flow velocity vector $\boldsymbol{v}_{\text{free}}$, expressed as:
$\boldsymbol{f}_{i}={{m}_{i}}\left(\boldsymbol{v}_{\text{free}}-\boldsymbol{v}_{i}\right)/\tau$ (2)
Here, $\tau$ represents the relaxation time, and $m_{i}$ denotes the mass of pedestrian $i$, $m_{i}/\tau$ can be regarded as a fixed proportional coefficient.
The interaction force between a pedestrian and the nearest pedestrian is represented by the repulsive force, governed by a power-law decay function with respect to distance, expanded as follows:
${\boldsymbol{f}_{ij}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{ij}}-\left\|{\boldsymbol{d}_{ij}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{2}}\left\|{\boldsymbol{d}_{ij}}\right\|^{\eta}\left(1+cos\theta\right)\right){\boldsymbol{n}_{ij}}$ (3)
Here, $c_{1}$, $c_{2}$, $\lambda_{1}$, $\alpha$ (not a necessary coefficient), and $\eta$ are proportional coefficients.The angle $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\boldsymbol{v}_{ij},\boldsymbol{d}_{ij}\right)\in\left[0,\pi\right]$. The symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\boldsymbol{v}_{ij}=\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$, and $\boldsymbol{d}_{ij}=\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$, ${\boldsymbol{n}_{ij}}=-{\boldsymbol{d}_{ij}}/\left\|{{\boldsymbol{d}_{ij}}}\right\|$.
The first term in the equation represents the contact force (elastic volume exclusion principle, exponentially decaying with distance), while the second term represents the non-contact force (repulsive effect, decaying with a power-law of distance). Given that, in low-density scenarios, the movement of pedestrian is also influenced by nearest neighbors from the far side (Murakami et al., 2021), a power-law function was employed rather than an exponential function for the calculation.
The interaction force between pedestrians and global obstacles is composed of two exponentially decaying terms with respect to distance, expanded as:
${\boldsymbol{f}_{iw}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{3}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{2}}}}\right){\boldsymbol{n}_{iw}}$ (4)
Here, ${{c}_{3}}$ and ${{\lambda}_{2}}\left({{\lambda}_{2}}\gg{{\lambda}_{1}}\right)$ represent proportionality coefficients, where $\left\|{\boldsymbol{d}_{iw}}\right\|$ denotes the perpendicular distance between pedestrian $i$ and the wall, and ${\boldsymbol{n}_{iw}}$ represents the standard vector opposite to the perpendicular line between pedestrian $i$ and the wall, ${\boldsymbol{n}_{iw}}=-{\boldsymbol{d}_{iw}}/\left\|{{\boldsymbol{d}_{iw}}}\right\|$.
The first term in the equation models the elastic volume exclusion principle exerted by the wall on the pedestrian under contact conditions, while the second term simulates the repulsive force of the wall on the pedestrian under non-contact conditions. If the virtual repulsive effect of the wall on the pedestrian is not required, the second term can be omitted. In this case, the wall can be considered a stationary obstacle composed of pedestrians.
3.2 Modeling Anticipation and Reaction Mechanisms via Cosine Functions
In this context, the anticipation mechanism of pedestrians is referred to as a cooperative game process among conflicting pedestrians (inducing phase separation). Conversely, the reaction mechanism is described as the convergence process among co-directional pedestrians (inducing aggregation).
The nearest-neighbor interaction force $\boldsymbol{f}_{ij}$ is adjusted nonlinearly using a cosine function of the angle $\theta$. When $\theta\to 0$, it indicates a potential collision trend, and $\boldsymbol{f}_{ij}$ increases. When $\theta\to\pi$, it indicates a separation trend between pedestrians, and $\boldsymbol{f}_{ij}$ decreases. With the introduction of the cosine function, the model approximately simulates the phase separation phenomenon between opposing pedestrians.
Figure 3: Schematic of nearest neighbor interaction: (a) Reciprocal interaction (anticipation mechanism), (b) Non-reciprocal interaction (reaction mechanism).
Phase separation of conflicting flows
Fig. 3(a) shows two agents, $i$ and $j$, in a cross-flow scenario with a potential collision, subject to mutually repulsive forces. Since agent $j$ is closer to the conflict point, the velocity reduction due to the repulsive force is smaller $\left(\left\|\boldsymbol{f}_{ji,y}\right\|<\left\|\boldsymbol{f}_{ij,x}\right\|\right)$. This trend is gradually amplified with the update of their relative positions, ultimately leading to agent $i$ yielding to agent $j$ (cooperative game). Strictly speaking, if the agents’ radii are ignored, non-cooperative games occur only when $\theta=0$.
Aggregation of co-directional flows
Fig. 3(b) illustrates that in unidirectional flow, agent $i$ is subjected to the repulsive force from agent $j$. If $\left\|\boldsymbol{v}_{j}\right\|>\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ increases, and $\boldsymbol{v}_{i}$ decreases. If $\left\|\boldsymbol{v}_{j}\right\|<\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ decreases, and $\boldsymbol{v}_{i}$ increases (considering the self-driving force). This leads to the aggregation of pedestrians in the same direction. Meanwhile, it satisfies the speed-density relationship under the constraints of the fundamental diagram.
4 Analytical Analysis
In the model, the force term is constrained by the exponential decay function and the power law decay function based on distance. In this section, a sensitivity analysis of the proportional coefficients governing the force term is performed.
4.1 Exponential Decay Term of Force
Fig. 4 illustrates the influence of the proportional coefficients $\lambda$ and $\alpha$ on the force term. A smaller $\lambda$ indicates that agents approach a rigid particle core, which may lead to a steep increase in interaction forces between pedestrians under high-density conditions, resulting in unrealistic instantaneous speeds. A larger value of $\lambda$ is indicative of pedestrian behavior resembling a collision with an elastic body; however, an excessively large $\lambda$ may result in an unrealistic overlap of body regions. The proportional coefficient $\alpha$ is used to adjust pedestrians’ dominance over nearby space. For a given decay coefficient ($\lambda$), increasing $\alpha$ enhances a pedestrian’s control over the neighboring area. Introducing $\alpha$ aims to simplify the setting of the coefficient $\lambda$. For an appropriate value of $\lambda$, the coefficient $\alpha$ can be removed from the model.
Figure 4: Sensitivity analysis of the nonlinear proportionality constants $\lambda$ (a) and $\alpha$ (b).
4.2 Power Law Decay Term of Force
Fig. 5 presents the influence of the proportional coefficient $\eta$ and angle $\theta$ on the power law decay force term. An increase in $\eta$ causes an increase in the decay rate of the force term. The parameter $\theta$ ensures that, for any given coefficient, when $\left\|{\boldsymbol{d}_{ij}}\right\|>{r}_{ij}$, ${f}_{ij}\in\left[0,2\right]$.
Figure 5: Sensitivity analysis of the nonlinear proportionality constant $\eta$ (a) and the angle $\theta$ (b).
The symbols setting are shown in Tab. 1.
Table 1: Symbol Descriptions and Reference Values
Symbol Description Reference value
$m$ Mass 60 kg
$v_{\rm{free}}$ Free-flow speed 1.2 m/s
$r$ Radius 0.2 m
$\tau$ Relaxation time 1 s
$c_{1},c_{2},c_{3}$ Proportionality coefficients (linear) \
$\lambda_{1},\lambda_{2},\alpha,\eta$ Proportionality coefficients (nonlinear) \
5 Conclusions
In this paper, we proposed a force-based general model: CosForce. In the model, we employed forces constrained by cosine functions to model pedestrians’ Anticipation mechanism (separation between conflicting pedestrians) and reaction mechanism (aggregation of co-directional flow). Despite its simplicity, this model demonstrates unexpectedly performance, capable of simulating diverse pedestrian self-organizing phenomena, such as lane formation, stripe formation, and cross-channel formation.
Currently, the quantitative evaluation of the model is ongoing. Demonstration videos and codes are available in the Data Availability section. The videos present rough simulation results of the model.
Data Availability
The codes and video of CosForce model can be found here: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link (Google Drive).
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 52072286, 71871189, 51604204), and the Fundamental Research Funds for the Central Universities (Grant No. 2022IVA108).